OpenStudy (fools101):

I need help! please Solve the following equation for Z 2z/7 = 14

OpenStudy (fools101):

Sorry doing Similarity*

OpenStudy (anonymous):

Multiply both sides by 7.

OpenStudy (fools101):

so i'm going to do 7*2 and 7*14 ????

OpenStudy (anonymous):

2z/7=14 2z=98 z=98/2 z=49

OpenStudy (fools101):

okay how do i do it to a / <---- ?

OpenStudy (fools101):

while i knew that sorry i should ask how to make in to / <--- that

OpenStudy (anonymous):

14*7/2

OpenStudy (anonymous):

2z/7=14 2z/7(7)=14(7) 2z=98 (The above, I did a shortcut. This one's the step that explains how i got 2z=98)

OpenStudy (anonymous):

I'm sorry, I do not follow?

OpenStudy (fools101):

OMG i'm so stupid soooo SORRY i was look at the wrong one thank u LMAO!! SEE i'm to tired haha been doing this all day!! >_< that embarrassing sorry

OpenStudy (anonymous):

No problem. You're welcome. I didn't get how solving for Z was a similarity problem. :)

OpenStudy (fools101):

lol that for sure make u think What the hell!! hahah thanks!!

OpenStudy (anonymous):

;) you're welcome!